OpenStudy (anonymous):

the curve x^2+4y^2=7+3xy a. show that there is a point p with the x coordinate 3 at which the line tangent to the curve at p is horizontal. find the y-coordinate of p. b. find the value of d^2y/dx^2 at the point p found in a. Does the curve have a local min, max, or neither at point p. please help >~<

OpenStudy (anonymous):

\[ x^2+4y^2=7+3xy\] \[2x +8yy'=3y+3xy'\] solve for \[y'\] set equal zero and see that x = 3 is a solution

OpenStudy (anonymous):

so you get y=2?